OpenStudy (anonymous):

If F'(x) = (1+t^3)^0.5 and F(1) = 5, then F(3) = A. 1.230 B. 3.585 C. 6.230 D. 8.535 E. 11.230

OpenStudy (anonymous):

\[ F(3)-F(1) = \int_1^3F'(x)~dx \]

OpenStudy (anonymous):

This means: \[ F(3) - 5 = \int_1^3 (1+t^3)^{0.5}~dx \]Though I have a hunch they meant to say \(F'(t) = (1+t^3)^{0.5}\)

OpenStudy (anonymous):

\[ F(3) = \int_1^3\sqrt{1+t^3}~dt + 5 \]

OpenStudy (anonymous):

Thank you!